Find (a) the current in each resistor I1= I2= I3=

Find

(a) the current in each resistor

I1=

I2=

I3=

(b) the power delivered to each resistor

P1=

P2=

P3=
Answer #1

let ia ,ibare two currents in two loops in clockwise direction.

apply kvl in first loop

-28ia-12ia+12ib-14+26=0

-40ia+12ib=-12..(1)

apply kvl in second loop

-12ib+12ia-8ib+14=0

12ia-20ib=-14.(2)

solve (1) and (2) equations

we get

ia=0.622 A

ib=1.073 A
I1=ia=0.622 A

i2=ib-ia=0.451 A

i3=ib=1.073A

power delivered to

R1=P1=i1^2*R1= 10.83 W

P2=i2^R2=2.44 W

P3=9.21 W
Answer #2

(a) the current in each resistor

I1=0.622

I2=0.451

I3=1.073

(b) the power delivered to each resistor

P1=10.8 W

P2=2.44 W

P3=9.2 w